# Action Spotting and Precise Event Detection in Sports: Datasets, Methods, and Challenges

### v1

@@ Abstract
Video event detection has become an essential component of sports analytics, enabling automated identification of key moments and enhancing performance analysis, viewer engagement, and broadcast efficiency. Recent advancements in deep learning, particularly Convolutional Neural Networks (CNNs) and Transformers, have significantly improved accuracy and efficiency in Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES). This survey provides a comprehensive overview of these three key tasks, emphasizing their differences, applications, and the evolution of methodological approaches. We thoroughly review and categorize existing datasets and evaluation metrics specifically tailored for sports contexts, highlighting the strengths and limitations of each. Furthermore, we analyze state-of-the-art techniques, including multi-modal approaches that integrate audio and visual information, methods utilizing self-supervised learning and knowledge distillation, and approaches aimed at generalizing across multiple sports. Finally, we discuss critical open challenges and outline promising research directions toward developing more generalized, efficient, and robust event detection frameworks applicable to diverse sports. This survey serves as a foundation for future research on efficient, generalizable, and multi-modal sports event detection.
@@ 1 Introduction
Sports represent one of the largest global markets, projected to reach 599.9 billion US dollars by 2025 and 826 billion US dollars by 2030, with a compound annual growth rate of 6.6% [1]. Beyond its economic impact across industries such as media, marketing, and apparel, sports fundamentally center around athletic performance, where optimizing player efficiency, refining game strategies, and enhancing fan engagement are critical.
@@
The rise of sports analytics—the systematic collection, processing, and analysis of performance data—has enabled data-driven decision-making, leading to fundamental shifts in strategy. For instance, basketball has seen an increased reliance on three-point shooting, guided by predictive models that estimate expected points from various court locations [2].
@@
Within this context, video event detection has emerged as a fundamental yet challenging task in sports analytics. Accurate identification of key moments—such as corner kicks in soccer, rally conclusions in racket sports, or scoring events across disciplines—provides critical insights for coaches and athletes, supporting more effective performance analysis and tactical decision-making. Moreover, event detection benefits downstream applications by filtering non-playing segments, optimizing computational resources for subsequent tasks such as object tracking, and enabling automated highlight generation for commercial broadcasting and fan engagement.
@@
Computer Vision (CV) has played a pivotal role in advancing sports video analysis, enabling automated player tracking, action recognition, tactical analysis, injury prevention, and officiating through advanced visual data processing [3, 4, 5, 6, 7, 8]. While traditional CV methods laid the foundation for many applications [5, 9], they heavily relied on handcrafted features and struggled with dynamic environments, occlusions, and real-time constraints.
Recent breakthroughs in deep learning (DL), particularly through Convolutional Neural Networks (CNNs) and Transformer architectures, have significantly enhanced the accuracy, efficiency, and scalability of sports video analysis tasks [5]. These modern techniques now power real-time object detection, robust player tracking, pose estimation, and fine-grained event recognition, fundamentally reshaping the landscape of sports analytics.
This survey provides a comprehensive analysis of DL-based CV techniques for video event detection in sports—a field that has seen remarkable advancements with the emergence of modern neural architectures and the formulation of distinct problem settings, notably Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES).
@@
Table 1: Comparison of Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES).
@@
Previous studies, such as by Ghosh et al. [5], broadly survey AI applications in sports analytics but do not specifically address DL-based CV methods. Similarly, Thomas et al. [9] focus primarily on traditional CV approaches designed for multi-camera systems. In contrast, our survey specifically targets deep learning-based approaches—including recent advances in CNNs and Vision Transformers—for event detection tasks within monocular video contexts, enhancing its relevance for real-world deployment.
Other comprehensive surveys by Naik et al. [3], Zhao et al. [4], Kamble et al. [10], and Wu et al. [6] extensively review computer vision methods across various sports analytics tasks, including object tracking and action recognition. However, they do not specifically address video event detection. The most closely related work by Hu et al. [11] provides an extensive overview of TAL methods, but it does not cover the increasingly critical tasks of AS and PES, nor is it specifically focused on the sports domain.
@@
  • Explicit Definition and Analysis of AS and PES: We clearly define, differentiate, and critically review AS and PES tasks, highlighting their significance and unique challenges within sports video analytics.
  • Comprehensive Methodological Survey: We systematically review DL-based approaches for sports event detection, including techniques based on convolutional networks, transformer architectures, and multi-modal fusion.
  • Evaluation Metrics and Datasets Review: We provide a thorough summary of evaluation protocols and benchmark datasets relevant to TAL, AS, and PES tasks.
  • Identification of Open Challenges: We discuss key limitations of existing methods, including issues related to temporal precision, multi-sport generalizability, and data efficiency, and outline promising research directions.
@@
Figure 1: Example of Temporal Action Localization: in tennis, the time interval of the whole serve action has been labeled (blue bar).
@@
2 Taxonomy of Sports Event Detection
In the domain of sports video analysis, three primary tasks relevant to event detection have been defined: TAL, AS, and PES. While these tasks share certain similarities, each addresses distinct objectives and annotation requirements, as summarized in Table 1. The key differences among these tasks are discussed in detail in the following subsections.
2.1 Temporal Action Localization
Temporal Action Proposal Generation (TAPG) aims to identify temporal segments within videos that are likely to contain actions or events, without necessarily assigning specific action labels. Building upon TAPG, Temporal Action Localization (TAL)—also referred to as Temporal Action Detection (TAD)—extends the task by not only detecting these temporal segments but also classifying the actions occurring within them [12, 13, 11], as illustrated in Figure 1.
Among these tasks, TAPG is generally regarded as the most challenging, as accurately estimating the temporal boundaries of actions is inherently difficult. In contrast, once reliable proposals are available, action classification models tend to achieve relatively high accuracy [11]. This makes TAL particularly effective for recognizing everyday actions that unfold over extended periods.
@@ 2.2 Action Spotting
Introduced by Giancola et al. in 2018 [14], AS—also referred to as Event Spotting—addresses the ambiguity inherent in defining precise temporal boundaries for actions. Several factors contribute to this challenge: some actions occur extremely rapidly, others have continuous and indistinct endpoints, and many actions may overlap or conflict with one another. These ambiguities are especially pronounced in sports scenarios. For instance, in soccer, accurately determining the start and end times of events such as ”scoring a goal” or ”passing” is often nontrivial, due to the continuous nature of gameplay and the lack of sharply defined action boundaries.
To address these challenges, AS proposes an alternative to TAL by representing each action with a single keyframe rather than a temporal interval. While both TAL and AS aim to temporally locate actions within untrimmed videos, they differ fundamentally in annotation granularity. TAL annotates actions with explicit start and end times, making it suitable for longer-duration activities. In contrast, AS captures brief and rapidly occurring events using a single timestamp, thereby simplifying annotation efforts and mitigating ambiguities associated with temporal boundaries.
@@ 2.3 Precise Event Spotting
While AS facilitates coarser detection by tolerating errors within relatively broad temporal windows (typically between 5 and 60 seconds), certain applications in sports demand much higher temporal precision. Hong et al. [15] highlighted that in sports such as tennis [16] and figure skating [17], even a margin of error of 1–2 frames can result in missing critical events, such as the precise moment of ball contact in tennis.
Similarly, consultations with Table Tennis Australia confirmed that frame-level precision is critical in table tennis, where even minor inaccuracies can significantly impact the determination of ball bounce locations, given the rapid motion of the ball across frames in professional-level matches, as illustrated in Figure 2.
To meet these demands, PES was proposed as an extension of AS. PES requires extremely strict detection thresholds, typically constrained within 1–5 frames (approximately 33–200 milliseconds at standard frame rates of 25–30 FPS). PES thus ensures highly accurate event localization at the frame level, addressing the critical need for fine-grained temporal precision in professional sports analysis.
@@ 3 Video Event Detection
Generally, a video event detection model do this task in two step:
  • Temporal Segmentation: This involves identifying the timing of an event, either by locating its start and end timestamps within a continuous video (as in Temporal Action Localization, TAL) or by detecting a single representative keyframe (as in Action Spotting, AS, and Precise Event Spotting, PES)
  • Event Classification: Once an event is segmented, the model classifies it into predefined categories, such as team events (e.g., goals, time-outs) or individual actions (e.g., shots, fouls, serves).
In this section, we dive into the review of video event detection.
3.1 Temporal Action Localization
Current methods for TAPG typically fall into two main categories: Global-to-Local and Local-to-Global approaches. Global-to-Local methods generate action proposals through sliding windows or pre-defined anchor structures, segmenting videos into fixed-length or hierarchically structured intervals to localize actions. Local-to-Global methods independently evaluate each frame to predict the likelihood of action start, action end, and overall actionness. Proposals are subsequently created by connecting high-confidence start and end frames, followed by a confidence-based evaluation to confirm whether the proposal indeed contains an action. The difference between these two approaches is shown in figure 3. The subsequent sections will explore these two approaches comprehensively, discussing their strengths, limitations, and practical applications in sports event detection.
@@ 3.1 Temporal Action Localization
To address these limitations, Complementary Temporal Action Proposal (CTAP) was introduced by J. Gao et al. [22], leveraging the strengths of both actionness-based proposals and sliding-window proposals. CTAP first generates both types of proposals and then applies a proposal complementary filter, which collects missing correct proposals that TAG alone might overlook by retrieving them from the sliding-window proposals. In this stage, a binary classifier is used, taking a sequence of unit features as input and outputting a probability score that indicates whether a proposal can be accurately detected by the unit-level actionness model and TAG. Finally, a Temporal Convolutional Adjustment and Ranking Network is employed to rank proposals and refine their boundaries. Experimental results on THUMOS-14 and ActivityNet v1.3 demonstrate that CTAP outperforms TURN in AR, highlighting its effectiveness in improving proposal quality. However, the combination of multiple proposal generation methods can impact computational efficiency, potentially increasing processing time.
Figure 4: Examples of events defined in the context of soccer. From top to bottom: Goal—the instant the ball crosses the goal line; Substitution—the moment a player enters the field; Card—the instant the referee shows a card to a player. All examples are from SoccerNet [14].

### v2

@@ Abstract
Video event detection is central to modern sports analytics, enabling automated understanding of key moments for performance evaluation, content creation, and tactical feedback. While deep learning has significantly advanced tasks like Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES), existing surveys often overlook the fine-grained temporal demands and domain-specific challenges posed by sports. This survey first provides a clear conceptual distinction between TAL, AS, and PES, then introduces a methods-based taxonomy covering recent deep learning approaches for AS and PES, including feature-based pipelines, end-to-end architectures, and multimodal strategies. We further review benchmark datasets and evaluation protocols, identifying critical limitations such as reliance on broadcast-quality footage and lenient multi-label metrics that hinder real-world deployment. Finally, we outline open challenges and future directions toward more temporally precise, generalizable, and practical event spotting in sports video analysis.
@@ 1 Introduction
Sports represent one of the largest global markets, projected to reach 599.9 billion US dollars by 2025 and 826 billion US dollars by 2030, with a compound annual growth rate of 6.6% [1]. Beyond its economic impact in industries such as media, marketing, and apparel, sports are fundamentally focused on athletic performance, where optimising player efficiency, refining game strategies, and enhancing fan engagement are critical. The rise of sports analytics, the systematic collection, processing, and analysis of performance data, has enabled data-driven decision-making, leading to fundamental changes in strategy. For instance, basketball has seen an increased reliance on three-point shooting, guided by predictive models that estimate expected points from various court locations [2].
@@
Within this context, video event detection has emerged as a fundamental yet challenging task in sports analytics. Accurate identification of key moments, such as corner kicks in soccer, rally conclusions in racket sports, or scoring events across disciplines, provides critical insight to coaches and athletes, supporting more effective performance analysis and tactical decision-making. Moreover, event detection benefits downstream applications by filtering non-playing segments, optimising computational resources for subsequent tasks such as object tracking, and enabling automated highlight generation for commercial broadcasting and fan engagement.
@@
Computer Vision (CV) has played a pivotal role in advancing sports video analysis, enabling automated player tracking, action recognition, tactical analysis, injury prevention and officiating through advanced visual data processing [3, 4, 5, 6, 7, 8]. While traditional CV methods laid the foundation for many applications [5, 9], they heavily relied on handcrafted features and struggled with dynamic environments, occlusions, and real-time constraints. Recent breakthroughs in Deep Learning (DL), particularly through Convolutional Neural Networks (CNNs) and Transformer architectures, have significantly enhanced the accuracy, efficiency, and scalability of sports video analysis tasks [5]. These modern techniques now power real-time object detection, robust player tracking such as [10, 11, 12], pose estimation [13], and fine-grained event recognition [14, 15], fundamentally reshaping the landscape of sports analytics.
Figure 1: Annual publication count from 2010 to 2024 based on a Scopus keyword search for ”sports” AND ”deep learning”.
To further illustrate the growing interest in this domain, we compiled a publication trend analysis using Scopus, based on keyword searches related to ”sports” and ”deep learning.” The resulting bar chart (Figure 1) shows a clear year-on-year increase in the number of relevant publications from 2010 to 2024, reinforcing the expanding research momentum in this field.
As video event detection in sports continues to evolve with the rise of DL-based CV techniques, a systematic literature review is essential to consolidate recent progress, clarify emerging task definitions—such as Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES)—and identify current methodological gaps and research opportunities.
@@
Table 1: Comparison of Temporal Action Localisation (TAL), Action Spotting (AS), and Precise Event Spotting (PES).
@@
Previous studies, such as those by Ghosh et al. [5], broadly survey AI applications in sports analytics but do not specifically address DL-based CV methods. Similarly, Thomas et al. [9] focus primarily on traditional CV approaches designed for multi-camera systems. In contrast, our survey specifically targets deep learning-based approaches—including recent advances in CNNs and Vision Transformers—for event detection tasks within monocular video contexts, enhancing its relevance for real-world deployment. Other comprehensive surveys by Naik et al. [3], Zhao et al. [4], Kamble et al. [16], and Wu et al. [6] extensively review CV methods across various sports analytics tasks, including object tracking and action recognition. However, they do not specifically address video event detection. The most closely related work by Hu et al. [17] provides an extensive overview of TAL methods, but it does not cover the increasingly critical tasks of AS and PES, nor is it specifically focused on the sports domain.
@@
  • Task-Oriented Taxonomy and Analysis: We introduce a structured taxonomy that formally defines and distinguishes AS and PES, positioning them within the broader landscape of video event detection. This taxonomy emphasizes the unique temporal and semantic challenges of sports analytics.
  • In-Depth Methodological Survey: We comprehensively review deep learning-based approaches for AS and PES, spanning feature-based pipelines, end-to-end architectures, and multi-modal frameworks leveraging convolutional, recurrent, and transformer-based designs.
  • Evaluation Protocols and Dataset Landscape: We summarize key benchmark datasets and evaluation metrics used across TAL, AS, and PES tasks, and critically examine their suitability for real-world deployment, including limitations in confidence thresholding and multi-label predictions.
  • Critical Insights and Future Directions: We identify open challenges such as poor generalization across diverse sports, limited supervision strategies, and unrealistic evaluation schemes. We propose actionable directions to guide the development of more robust and deployable spotting models.
@@
The remainder of the paper is organized as follows. Section 2 provides clear definitions of the various tasks involved in sports video event detection. In Section 3, we review existing methodologies developed for this domain. Section 4 and 5 introduce relevant evaluation metrics and datasets commonly used for sports event detection. Section 6 outlines key challenges and future research directions. Finally, Section 7 concludes the paper with a summary of insights and findings.
Figure 2: Example of Temporal Action Localization: in tennis, the time interval of the whole serve action has been labelled (blue bar).
@@
2 Sports Event Detection
In sports video analysis, three key tasks have emerged for temporal event detection: Temporal Action Localisation (TAL), AS, and PES. While related, they differ in output format, annotation granularity, and application focus (see Table 1).
2.1 Temporal Action Localisation
TAL—also referred to as Temporal Action Detection (TAD)—aims to detect and classify action segments within untrimmed videos. It builds on Temporal Action Proposal Generation (TAPG), which identifies candidate segments likely to contain actions, without assigning labels [18, 19, 17].
Although originally designed for general activity understanding, many TAL methods have been adapted to sports due to their strength in modeling temporal structure. TAPG is especially challenging because it requires precise boundary prediction, while classification within those segments is typically more accurate once good proposals are established [17].
@@ 2.2 Action Spotting
AS, introduced by Giancola et al. [20], addresses the difficulty of defining precise action boundaries in sports, where events are fast, overlapping, or continuous. Instead of annotating start and end times, AS uses a single timestamp to represent an event, simplifying annotation and avoiding ambiguity.
While both TAL and AS seek to locate events temporally, AS trades interval-level precision for a more efficient and scalable annotation strategy, better suited for rapid and ambiguous sports actions.
@@ 2.3 Precise Event Spotting
PES extends AS by enforcing strict frame-level precision. In sports like tennis [21] and figure skating [22], even a 1–2 frame error can miss critical events. Similarly, feedback from Table Tennis Australia confirms that pinpoint accuracy is essential for detecting ball contact or bounce points, where the ball may travel across frames in under 100 milliseconds.
PES thus requires event detection within tight tolerance windows (typically 1–5 frames), making it ideal for high-precision, frame-accurate tasks in elite sports contexts.
@@ 3 Video Event Detection
In this section, we provide a comprehensive overview of video event detection methods, covering both general-purpose approaches and those specifically designed for sports. We begin by reviewing foundational techniques originally developed for generic activity detection, many of which have influenced early progress in sports video analysis. We then shift our focus to sports-specific methodologies, including the recent emergence of tasks such as AS and PES, which address the unique challenges of detecting fine-grained events in fast-paced, ambiguous sports scenarios.
3.1 General Video Event Detection
Current methods for TAL typically fall into two main categories: Global-to-Local and Local-to-Global approaches. Global-to-Local methods generate action proposals through sliding windows or pre-defined anchor structures, segmenting videos into fixed-length or hierarchically structured intervals to localise actions. Local-to-Global methods independently evaluate each frame to predict the likelihood of action start, action end, and overall actionness. Proposals are subsequently created by connecting high-confidence start and end frames, followed by a confidence-based evaluation to confirm whether the proposal contains an action. The difference between these two approaches is shown in Figure 4. The subsequent sections will explore these two approaches comprehensively, discussing their strengths, limitations, and practical applications in sports event detection.
@@ 3.1.1 Global to Local
To address these limitations, the Complementary Temporal Action Proposal (CTAP) was introduced by J. Gao et al. [27], leveraging the strengths of both actionness-based proposals and sliding-window proposals. CTAP first generates both types of proposals and then applies a proposal complementary filter, which collects missing correct proposals that TAG alone might overlook by retrieving them from the sliding-window proposals. In this stage, a binary classifier is used, taking a sequence of unit features as input and outputting a probability score that indicates whether a proposal can be accurately detected by the unit-level actionness model and TAG. Finally, a Temporal Convolutional Adjustment and Ranking Network is employed to rank proposals and refine their boundaries. Experimental results on THUMOS-14 and ActivityNet v1.3 demonstrate that CTAP outperforms TURN in AR, highlighting its effectiveness in improving proposal quality. However, the combination of multiple proposal generation methods can impact computational efficiency, potentially increasing processing time.
GTL methods follow a two-stage paradigm similar to R-CNN: they first generate candidate proposals and then classify them. This approach has shown strong performance in early temporal action localization work. However, a key limitation lies in the fixed-length nature of the proposals, which does not align well with the highly variable durations of action instances across different videos. As a result, these methods often rely on dense sliding windows to ensure sufficient coverage, leading to high computational overhead. Furthermore, while classification accuracy is generally strong, GTL approaches often struggle to produce precise temporal boundaries, limiting their effectiveness for fine-grained event localization [17].
Figure 5: Examples of events defined in the context of soccer. From top to bottom: Card—the referee shows a card to a player; Substitution—the moment a player enters the field; and Goal—the ball crosses the goal line. All examples are from SoccerNet [20].
3.1.2 Local to Global
@@
Compared to GTL, LTG methods address the limitation of fixed proposals by enabling flexible boundary generation through frame-level classification. This allows for more adaptive localization of action segments. However, LTG approaches are often sensitive to fluctuations in boundary predictions, which can result in unstable or fragmented proposals [17].
Figure 6: Illustration of three representative modeling strategies for AS and PES tasks. Pooling-Based Models perform temporal aggregation on extracted features; Encoder-Based Models use temporal encoders (e.g., 1D/3D CNNs, Transformers) to capture sequential dependencies; and Frame-Discriminative Models modify the backbone to integrate frame-level temporal reasoning directly for improved spotting accuracy.
Figure 7: Generalized architecture for AS and PES models. Most approaches follow a four-stage modular pipeline: input sampling (frame-wise or chunk-based), feature extraction (via 2D/3D CNNs or transformers), temporal modeling (e.g., 1D CNNs, RNNs, or self-attention), and classification. This structure underlies a variety of modeling strategies across the AS and PES literature.
3.2 Sports Video Event Detection
While general video event detection methods laid the foundation for sports analytics, the unique demands of sports—such as rapid actions, ambiguous event boundaries, and the need for frame-level precision—have driven the development of specialized approaches. TAL methods, originally designed for general activity understanding, often fall short in these contexts due to their reliance on coarse temporal intervals.
To address these limitations, two sports-specific tasks have emerged: AS and PES. Unlike TAL, which focuses on identifying action segments, AS and PES aim to detect the exact moment an event occurs. This shift from segment-based to keyframe-based detection better reflects the temporal granularity required in sports settings.
Although AS and PES differ in evaluation metrics and frame-level precision requirements, many detection models are applicable to both. We categorize existing approaches according to their underlying architectural strategies:
  • Pooling-Based Methods aggregate features from pretrained models within a fixed temporal window using different pooling techniques.
  • Temporal Encoder-Based Methods employ sequence models (e.g., 1D/3D CNNs or Transformers) to model temporal dependencies across frames.
  • Frame-Discriminative Models focus on designing architectures to capture fine-grained frame-level differences, enabling more precise temporal localization tailored to the demands of spotting tasks.
  • Multi-Modal Fusion Methods incorporate non-visual modalities, such as audio, to enhance detection performance.
  • Data-Efficient Learning Approaches reduce reliance on large-scale annotations through active learning, self-supervised learning, or knowledge distillation.
Figure 6 presents a visual overview of three representative architectural strategies—Pooling-Based Models, Encoder-Based Models, and Frame-Discriminative Models—that form the core of most AS and PES approaches. Multi-modal methods typically adopt one of these architectures while incorporating additional input sources such as audio or text. In contrast, data-efficient learning approaches (e.g., self-supervised or active learning) represent a separate paradigm focused on training strategy rather than model structure, and are thus not easily categorized within the same architectural taxonomy.
While the taxonomy in Figure 6 organizes AS and PES methods by high-level architectural strategy, most models—regardless of category—follow a shared modular pipeline. This typically consists of input sampling, frame-level feature extraction, temporal modeling, and classification. Figure 7 illustrates this generalized structure, which underpins both frame-wise and chunk-based approaches, and helps contextualize where architectural variations occur within the overall flow.
3.2.1 Pooling Based Approaches
Pooling-Based Approaches aggregate features over fixed temporal windows using techniques like average pooling, NetVLAD, or its temporally-aware variant NetVLAD++. These models typically use 2D or 3D CNNs to extract frame-level features, followed by simple aggregation and classification.
One of the earliest works in this category is the SoccerNet baseline [20], which used average and max pooling on pre-extracted features from C3D [24], I3D [34], and ResNet [35]. Rongved et al. [36] later extended this with a 3D ResNet—adapted from [37]—using a 5-second sliding window to better capture temporal dynamics. NetVLAD++ [38] further improved upon these ideas by explicitly modeling temporal order in the aggregation process, enabling better contextual understanding of past and future frames.
Zhou et al. [39] advanced this line of work by fine-tuning multiple action recognition backbones—including TPN [40], GTA [41], VTN [42], irCSN [43], and I3D-Slow [44]—on soccer video snippets. The combined features, when processed by NetVLAD++ or Transformer-based encoders, achieved state-of-the-art performance on the SoccerNet benchmark.
Zhu et al. [45] introduced a more efficient approach by utilizing a single multi-scale Vision Transformer for feature extraction, followed by NetVLAD++ pooling. This model strikes a balance between temporal modeling capacity and computational efficiency, making it more suitable for deployment in resource-constrained settings.
To better leverage the structure of events over time, Cioppa et al. [46] proposed a context-aware loss function that dynamically weights frames based on their proximity to annotated events. This smooth temporal weighting significantly improved focus on relevant cues, resulting in a 12.8% performance boost on SoccerNet-v1. However, its effectiveness diminishes in denser datasets like PES [47], which require higher frame-level precision.
3.2.2 Encoder-Based Methods
Encoder-Based Methods enhance feature exploitation pipelines by replacing pooling operations with sequence models that learn temporal dependencies directly from the input features. These models include 1D CNNs, 3D CNNs, and Transformers, which allow for richer contextual modeling.
To address variability in action duration, Vats et al. [48] proposed a multi-tower CNN architecture operating at multiple temporal resolutions. This multi-scale design captures both short-term and long-range dependencies, improving robustness across diverse sporting actions.
Tomei et al. [49] introduced RMS-Net, a hybrid model combining regression-based localization with a masking strategy inspired by object detection and self-supervised learning. This design promotes the learning of temporally structured signals, improving the precision of event boundaries.
SpotFormer [50] further demonstrates the strength of encoder-based models by fusing features from multiple pretrained backbones (e.g., VideoMAE [51], Swin Transformer [52]) and processing them with a Transformer spotting head. While highly accurate, this approach is computationally intensive and less suitable for real-time use.
To address computational constraints, Darwish et al. [53] proposed the Spatio-Temporal Encoder (STE), a lightweight model employing 1D convolutions and MLPs. Despite its simplicity, STE delivers competitive performance by efficiently capturing key temporal patterns.
Soares et al. [54] adapted an anchor-based detection framework—originally designed for TAL—to AS. Their model combines a 1D U-Net with a Transformer encoder, effectively capturing both local and long-range temporal dependencies, as reflected in their strong performance on SoccerNet Challenge [55].
Shi et al. [56] addressed variable-length events through a multi-scene encoding strategy, segmenting video chunks into variable-length subsets processed by separate Transformers. This enables the model to adaptively capture temporal context across varying durations.